How far must an object move in order for a 560 force to perform a 3.68 x104 J, if the angle is 60° ?


1
Expert's answer
2021-03-21T11:21:12-0400

"W=F\cdot s\cdot \cos\alpha\to s=\frac{W}{F\cdot\cos\alpha}=\frac{36800}{560\cdot\cos60°}=131.4\ (m)" . Answer